OpenStudy (anonymous):

What is the simplified form of the expression? 33/4 q/q^8

OpenStudy (unklerhaukus):

is that:\[\frac{33}4\times\frac q{q^8}\]?

OpenStudy (sepeario):

^^ probably.

OpenStudy (anonymous):